OpenStudy (anonymous):

Fill in the missing value. 1 km2 = ___________ dm2

mathslover (mathslover):

\( \begin{array}{} \cal{\large{\text{Hey! Good to see you on this site!}}}\\ \\ {\huge \mathtt{\color{blue}{\text{Welcome To OpenStudy!}}}} \\ \text{enjoy answering and asking questions here}\\ \cal{\large{\text{Tips:)}}} \\ \textbf{The people who have their name in purple are moderators}\\ \textbf{You have to obey their orders !}\\ \textbf{Want to join more groups ? click 'Find More Subjects'. }\\ \textbf{Have any doubt ! Ask us!}\\ \textbf{OS Code Of Conduct }\\ \textbf{openstudy.com/code-of-conduct}\\ \end{array}\)

mathslover (mathslover):

1 kilometer = 10 000 decimeters hence \[\huge{1km^2=(10000dm)^2}\]

OpenStudy (anonymous):

oh great

OpenStudy (anonymous):

:)

mathslover (mathslover):

can u solve this now ?

mathslover (mathslover):

ur welcome @HoneyDarlin

OpenStudy (anonymous):

100,000,000? I really suck at math sorry lol

mathslover (mathslover):

yes : u r right gr8 \[\huge{10000^2=100000000=10^8}\] u can include \(\huge{dm^2}\) there after \(\huge{10^8}\)